OpenStudy (anonymous):

What is the first step in solving the following inequality? A. Square both sides B. Divide both sides by 2 C. Add 7 to both sides. D. Divide both sides by 3.

OpenStudy (anonymous):

Whats your guess???

OpenStudy (anonymous):

and whats the inequality

OpenStudy (anonymous):



OpenStudy (ryan123345):

What did i tell you to do last time?

OpenStudy (anonymous):

I think its B

OpenStudy (ryan123345):

no.

OpenStudy (anonymous):

So i should sqaure both sides first then ?

OpenStudy (anonymous):

But i thought you had to get rid of the two first ?

OpenStudy (anonymous):

In your given inequality, \[2\sqrt{3x-7}\ge1\]I can say that whatever works best for you. If you think dividing the inequality to 2 is the first easier step, then make it the first step. Or, getting the square root of both sides are easier, then make it your first step. ==> To get the inequality in a simpler form, you are right. Divide both sides by 2\[\sqrt{3x-7}\ge1/2\] \[(\sqrt{3x-7})^2\ge(1/2)^2\] \[3x-7\ge\frac{1}{4}\] \[3x \ge\frac{1}{4}\] \[3x \ge \frac{ 7+4 }{ 28 }\] \[3x \ge \frac{ 11 }{ 28 }\]\[\frac{ 3x }{3 } \ge \frac{ 11 }{ 28*3 }\]\[x \ge \frac{ 11 }{ 84 }\]